OpenStudy (anonymous):

Convert each of the following equations from polar coordinates to rectangular coordinates. r^2sin(2theta)=4

OpenStudy (rsadhvika):

hint : \(\large \sin(2\theta) = 2\sin \theta \cos\theta\)

OpenStudy (anonymous):

do i divide both sides next?

OpenStudy (rsadhvika):

\[\large r^2\sin(2\theta) = r^2\cdot 2\sin\theta \cos\theta = 2(r\sin\theta)(r\cos\theta) = ?\]

OpenStudy (rsadhvika):

\(\large x = r\cos\theta\) \(\large y = r\sin\theta\) remember these ?

OpenStudy (anonymous):

yes

OpenStudy (rsadhvika):

just plug them in and you're done

OpenStudy (rsadhvika):

\[\large r^2\sin(2\theta) = r^2\cdot 2\sin\theta \cos\theta = 2(r\sin\theta)(r\cos\theta) = 2xy\]

OpenStudy (rsadhvika):

so the rectangular form would be : \[\large 2xy = 4\]

OpenStudy (rsadhvika):

which is same as \[\large xy = 2\]